OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, 6) and is perpendicular to the line x = 4. y = 4 y = 6 x = 6 x = 2

OpenStudy (texaschic101):

y = 6......because x = 4 is a vertical line (it goes up and down)...and if you want a perpendicular line that goes through 6 on the y axis....then it would be y = 6. I am not real good at explaining this stuff...

OpenStudy (texaschic101):

NNesha....can you explain this better ?

OpenStudy (texaschic101):

understand this....when you have y = a number, then it is a horizontal line (going across) and if you have x = a number, then it is a vertical line (going up and down)

OpenStudy (anonymous):

Its ok i kinda get it

OpenStudy (anonymous):

Ready for the next one? here you go

Nnesha (nnesha):

slope of vertical line is undefined doesn't matter what number we get at the numerator but denominator would always be zero so let's say slope of vertical line \[\large\rm m = \frac{ 3 }{ 0 }\] and slope of perpendicular will be negative reciprocal flip the fraction and change the sign \[\rm m= -\frac{0}{3}\] ^^^^^^^ slope of perpendicular line plug that into the original equation \[\large\rm y=mx+b\] m is slope and b is y-intercept substitute x and y order pair for (2,6) \[\rm 6=-\frac{0}{3}(2)+b\] 0 times anything = same thing so \[\rm b=\color{red}{6} \rightarrow y=\color{reD}{b} \rightarrow y=\color{ReD}{6}\]